# Quark-Gluon Plasma, Charmonia, Ultra-relativistic Heavy-Ion Collisions

### v1

@@ 5 Conclusions
We have investigated the production of charmonia in ultra-relativistic heavy-ion collisions using a previously constructed semi-classical transport approach that satisfies detailed balance and incorporates gradual quarkonium dissociation utilizing reaction rates based on in-medium binding energies and heavy-quark masses. Compared to our previous studies, notable updates include revised in-medium binding energies guided by recent $T$-matrix computations, state-of-the-art charm production cross sections from experiment and their $p_{\rm T}$ dependent shadowing. Our focus has been on charmonium kinetics in 5.02 TeV Pb-Pb collisions, and specifically on predictions for $\psi(2S)$ observables that recently became available at the LHC. An important role is being played by the mechanism of “sequential regeneration", where regeneration processes for the $\psi(2S)$ are operative at lower temperatures than for the $J/\psi$, with significant contributions also from the hadronic phase. While the total $J/\psi$ yield is close to its chemical equilibrium values in the QGP phase of central Pb-Pb collisions (around temperatures of $\sim$250 MeV), the $\psi(2S)$ yields chemically equilibrate later, at temperatures of $\sim$160 MeV. This delay has significant consequences for observables, most notably a $\psi(2S)/J/\psi$ ratio above the equilibrium values at any given temperature and a shift of the “flow bump" in the nuclear modification factor to higher momenta for $\psi(2S)$ than for $J/\psi$. The former has been confirmed by experiment, while the latter is a more subtle effect that the data are not (yet) sensitive to. Furthermore, we have also re-confirmed the rather negligible effect of gluo-dissociation on the reaction rates and highlighted limitations of our blastwave approximation for the $p_{\rm T}$ spectra of the regenerated charmonia, which leads to an overestimation of the low-$p_{\rm T}$ yields in peripheral Pb-Pb collisions. On the other hand, in central collisions the $p_{\rm T}$-dependent $R_{\rm AA}$’s align well with the ALICE data, indicating that the charm-quark spectra at low and intermediate momenta are near local thermal equilibrium. Improvements by implementing the full kinetics of charm-quark diffusion have already been worked out for specific cases (and enabled, e.g., a resolution of the so-called $J/\psi$ $v_{2}$ puzzle He et al. (2022)), but are still awaiting systematic applications to the full available data samples. Further objectives of future developments are the implementation of nonperturbative matrix elements for the quasifree processes, and a realistic implementation of quantum transport for charmonia Andronic et al. (2024) that, in particular, can cope with regeneration reactions in the presence of multiple charm-anticharm quark pairs. Work in all these directions is in progress.

### v2

@@ 5 Conclusions
We have investigated the production of charmonia in ultra-relativistic heavy-ion collisions using a previously constructed semi-classical transport approach that satisfies detailed balance and incorporates gradual quarkonium dissociation utilizing reaction rates based on in-medium binding energies and heavy-quark masses. Compared to our previous studies, notable updates include revised in-medium binding energies guided by recent $T$-matrix computations, state-of-the-art charm production cross sections from experiment and their $p_{\rm T}$ dependent shadowing. Our focus has been on charmonium kinetics in 5.02 TeV Pb-Pb collisions, and specifically on predictions for $\psi(2S)$ observables that recently became available at the LHC. An important role is being played by the mechanism of “sequential regeneration", where regeneration processes for the $\psi(2S)$ are operative at lower temperatures than for the $J/\psi$, with significant contributions also from the hadronic phase. While the total $J/\psi$ yield is close to its chemical equilibrium values in the QGP phase of central Pb-Pb collisions (around temperatures of $\sim$250 MeV), the $\psi(2S)$ yields chemically equilibrate later, at temperatures of $\sim$160 MeV. This delay has significant consequences for observables, most notably a $\psi(2S)/J/\psi$ ratio above the equilibrium values at any given temperature and a shift of the “flow bump" in the nuclear modification factor to higher momenta for $\psi(2S)$ than for $J/\psi$. The former has been confirmed by experiment, while the latter is a more subtle effect that the data are not (yet) sensitive to. Furthermore, we have also re-confirmed the rather negligible effect of gluo-dissociation on the reaction rates and highlighted limitations of our blastwave approximation for the $p_{\rm T}$ spectra of the regenerated charmonia, which leads to an overestimation of the low-$p_{\rm T}$ yields in peripheral Pb-Pb collisions. On the other hand, in central collisions the $p_{\rm T}$-dependent $R_{\rm AA}$’s align well with the ALICE data, indicating that the charm-quark spectra at low and intermediate momenta are near local thermal equilibrium. Improvements by implementing the full kinetics of charm-quark diffusion have already been worked out for specific cases (and enabled, e.g., a resolution of the so-called $J/\psi$ $v_{2}$ puzzle He:2021zej), but are still awaiting systematic applications to the full available data samples. Further objectives of future developments are the implementation of nonperturbative matrix elements for the quasifree processes, and a realistic implementation of quantum transport for charmonia Andronic:2024oxz that, in particular, can cope with regeneration reactions in the presence of multiple charm-anticharm quark pairs. Work in all these directions is in progress.